You wish to make a 0.253 M hydroiodic acid solution from a stock solution of 12.0 M hydroiodic acid. How much concentrated acid must you add to obtain a total volume of 50.0 mL of the dilute solution? in mL

In the laboratory you dilute 2.31 mL of a concentrated 6.00 M hydrobromic acid solution to a total volume of 100 mL. What is the concentration of the dilute solution?

You wish to make a 0.388 M hydrochloric acid solution from a stock solution of 3.00 M hydrochloric acid. How much concentrated acid must you add to obtain a total volume of 50.0 mL of the dilute solution? in mL
Answer #1

1)

use dilution formula

M1*V1 = M2*V2

1---> is for stock solution

2---> is for diluted solution

Given:

M1 = 12 M

M2 = 0.253 M

V2 = 50 mL

use:

M1*V1 = M2*V2

V1 = (M2 * V2) / M1

V1 = (0.253*50)/12

V1 = 1.0542 mL

Answer: 1.05 mL

2)

use dilution formula

M1*V1 = M2*V2

1---> is for stock solution

2---> is for diluted solution

Given:

M1 = 6 M

V1 = 2.31 mL

V2 = 100 mL

use:

M1*V1 = M2*V2

M2 = (M1*V1)/V2

M2 = (6*2.31)/100

M2 = 0.1386 M

Answer: 0.139 M

3)

use dilution formula

M1*V1 = M2*V2

1---> is for stock solution

2---> is for diluted solution

Given:

M1 = 3 M

M2 = 0.388 M

V2 = 50 mL

use:

M1*V1 = M2*V2

V1 = (M2 * V2) / M1

V1 = (0.388*50)/3

V1 = 6.47 mL

Answer: 6.47 mL